OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (7, 5) and (7, 11). is it (0,6)?

OpenStudy (lgbasallote):

\[\huge \left( \frac{7 + 7}{2}, \; \frac{5+11}{2} \right)\] is that what you did?

OpenStudy (lgbasallote):

that came from the midpoint formula \[\huge \left(\frac{x_1 + x_2}{2} ,\; \frac{y_1 + y_2}{2} \right)\]

OpenStudy (anonymous):

oh so its (7,8).

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

im thinking you did \[\huge \left(\frac{7-7}{2} , \; \frac{11-5}{2}\right)\] didnt you?

OpenStudy (lgbasallote):

oh wait you just did \[\huge (7-7, \; 11-5)?\]